OpenStudy (anonymous):

–7 + g = 13 A. –20 B. –6 C. 6 D. 20

OpenStudy (anonymous):

Similar to your other question, simply move (-7) to the other side to make 'g' your subject Note that 'signs' change as you move them across the equal sign :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so do you do 7 - (-7)

OpenStudy (anonymous):

@DelTaVsPi

OpenStudy (anonymous):

@Music_me

OpenStudy (anonymous):

Don't think of it that way, it is more confusing its more like: - 7 + g = 13 g = 13 + 7

OpenStudy (anonymous):

Simply more (-7) across the equal sign the make it positive :) So it becomes +7

OpenStudy (anonymous):

so 20? @DelTaVsPi

OpenStudy (anonymous):

That is correct :)

OpenStudy (anonymous):

If you need any further assistance, please don't mind mentioning me ^^

OpenStudy (anonymous):

k thankyouu

OpenStudy (anonymous):

No problem, anytime !

OpenStudy (anonymous):

im sorry but could you walk me through some more of these questions because i miss this lesson because i wasnt enrolled into this school on this semester that this lesson is on

OpenStudy (anonymous):

Ah okay sure no problem !

OpenStudy (anonymous):

Solve. h + (–16) = –9 A. –25 B. –7 C. 7 D. 25

OpenStudy (anonymous):

@DelTaVsPi

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Hey sorry for the wait : ( The answer to that is 7, why? because +(-16) = -16 So by moving across the equal sign, it becomes positive So its shows as -9 + 16 = 7

OpenStudy (anonymous):

thx to both of you

OpenStudy (anonymous):

Welcome ;)